OpenStudy (anonymous):

Can you please show me an example on how to make fractions into percentages?

OpenStudy (haleyelizabeth2017):

All you need to do is divide the numerator by the denominator. for example, 5/10 = \(5 \div 10\) =0.5*100=50%

OpenStudy (anonymous):

50/100 = 50 percent 50 divided by 100 = .5 = 50 %

OpenStudy (anonymous):

But what happens if the top number is higher than the bottom?

OpenStudy (anonymous):

@buyfn @AlexandervonHumboldt2

OpenStudy (phi):

6/5 becomes 1 and 1/5 or as a decimal, 1.2 changed to percent: multiply by 100 and tack on a % sign: 1) 1.2*100 = 120 2) tack on % sign: 120% so 6/5 is 120% 100% means an "entire whole" 120% means 100% (an entire whole) plus 20% more

Directrix (directrix):

Here is a visual. http://www.slideshare.net/tdimella/fractions-decimals-and-percents-13452101

OpenStudy (anonymous):

Thanks guys ^3^ love ya~